OpenStudy (anonymous):

divide (10-6i)/(2+3i) a.12/13+42/13i b.2/13-42/13i c.5-2i d.3

OpenStudy (anonymous):

\[\frac{10-6i}{2+3i}=\frac{10-6i}{2+3i}\times \frac{10-6i}{2-3i}\]

OpenStudy (anonymous):

Multiply the conjugate of the denominator with the numerator and the denominator

OpenStudy (anonymous):

denominator is \[2^2+3^2=4+9=13\]

OpenStudy (anonymous):

ooops typo my mistake sorry.

OpenStudy (anonymous):

2(5-3i) / (2 +3i) i don't know whats next

OpenStudy (anonymous):

\[\frac{10-6i}{2+3i}\times \frac{2-3i}{2-3i}\]

OpenStudy (anonymous):

denominator is 13 numerator is \[(10-6i)(2-3i)=2-42i\]

OpenStudy (anonymous):

so you get \[\frac{2-42i}{13}=\frac{2}{13}-\frac{42}{13}i\] in standard form

OpenStudy (radar):

38-41i ------- 13

OpenStudy (anonymous):

radar -30-12=-42 i think

OpenStudy (radar):

Yes, I see that I erred. Lost track of signs lol

OpenStudy (anonymous):

and \[-6 \times (-3)=+18\] making the real part 20-18=2

OpenStudy (radar):

I totally agree satellite, please excuse my mistake Adam12.

OpenStudy (radar):

Looks like choice B is the best one.